# VisionTS: Visual Masked Autoencoders Are Free-Lunch Zero-Shot Time Series Forecasters

### v1

@@ Abstract
Foundation models have emerged as a promising approach in time series forecasting (TSF). Existing approaches either fine-tune large language models (LLMs) or build large-scale time-series datasets to develop TSF foundation models. However, these methods face challenges due to the severe cross-domain gap or in-domain heterogeneity. In this paper, we explore a new road to building a TSF foundation model from rich and high-quality natural images, based on the intrinsic similarities between images and time series. To bridge the gap between the two domains, we reformulate the TSF task as an image reconstruction task, which is further processed by a visual masked autoencoder (MAE) self-supervised pre-trained on the ImageNet dataset. Surprisingly, without further adaptation in the time-series domain, the proposed VisionTS could achieve superior zero-shot forecasting performance compared to existing TSF foundation models. With minimal fine-tuning, VisionTS could further improve the forecasting and achieve state-of-the-art performance in most cases. These findings suggest that visual models could be a free lunch for TSF and highlight the potential for future cross-domain research between computer vision and TSF. Our code is publicly available at https://github.com/Keytoyze/VisionTS.
Figure 1: Long-term forecasting performance comparison. Our VisionTS, without any training on time series data, outperforms the largest foundation model MoiraiLarge in the zero-shot setting.
@@ 1 Introduction
We have recently witnessed two roads to building a TSF foundation model. The first tries to fine-tune the weights of large language models (LLMs) which have been pre-trained on text data for TSF tasks (i.e., text-based) (Zhou et al., 2023; Jin et al., 2024), based on the observation that LLMs and TSF models share a similar left-to-right forecasting paradigm. However, due to the significant gap between these two modalities, the effectiveness of such transferability between language and time series has recently been questioned by Tan et al. (2024).
@@
The second road focuses on constructing large-scale time-series datasets collected from diverse domains to train a TSF foundation model from scratch (i.e., TS-based) (Woo et al., 2024; Das et al., 2024). Nevertheless, unlike images or language with unified formats, time series data is highly heterogeneous in length, frequency, dimensionality, domains, and semantics, limiting the transferability between pre-training and downstream domains. Until recently, constructing a high-quality dataset remains challenging and is still in the early exploration stage.
@@
In this paper, we investigate a third road that is less explored yet promising: building TSF foundation models with pre-trained visual models. Our starting point is the intrinsic similarities between natural images and TS: 1 Similar modalities: Unlike discrete texts, both images and TS are continuous. 2 Similar origin: Both TS and images are observations of real-world physical systems, whereas languages are products of human cognitive processes. 3 Similar information density: Languages are human-generated signals with high semantic density, while images and TS data are natural signals with heavy redundancy (He et al., 2022). 4 Similar features: As shown in Fig. 2, images often display many features of real-world time series, which are rarely found in language data. These findings suggest that visual models could be easier to adapt to TSF tasks than language models. Based on these insights, we are motivated to answer the question: “Can a visual model pre-trained on images be a free-lunch foundation model for zero-shot time series forecasting?
@@
Figure 2: An image of the ImageNet dataset (Deng et al., 2009), in which the pixel arrays can display many well-known features of real-world time series, such as trend, seasonality, and stationarity (Qiu et al., 2024). By self-supervised pre-training on ImageNet, it is reasonable that a visual model could understand these features and exhibit a level of time-series forecasting ability.
@@
We focus on visual masked autoencoder (MAE) [^1] We use fonts to distinguish MAE (Masked Autoencoder) and MAE (Mean Absolute Error) in this paper., a popular CV foundation model (He et al., 2022) by self-supervised pre-training on ImageNet (Deng et al., 2009). Inspired by the well-known prompt technique in NLP (Schick & Schütze, 2021), we propose a simple method to reformulate TSF as a patch-level image reconstruction task to bridge the gap between pre-training and downstream tasks. Specifically, we transform 1D TS data into 2D matrices via segmentation. Then, we render the matrices into images and align the forecasting window with masked image patches. This method allows us to make zero-shot forecasting without further adaptation.
@@
We evaluate our proposed VisionTS on multiple TSF benchmarks, including long-term TSF datasets (Zhou et al., 2021; Wu et al., 2021) and Monash (Godahewa et al., 2021). As demonstrated in Fig. 1, without any further adaptation in the time-series domain, a vanilla MAE can surprisingly achieve a comparable performance or even outperform the state-of-the-art (SOTA) zero-shot TSF foundation models, including text-based and TS-based methods. By fine-tuning MAE on each downstream dataset for only one epoch, VisionTS can lead to a SOTA performance on most long-term TSF benchmarks. Our findings suggest that time series and natural images may be two sides of a coin, and visual models can be a free lunch for time series forecasting. We hope our findings can inspire future cross-domain research on CV and TSF.
@@
  • We explore a road to building a TSF foundation model from natural images, which is conceptually different from the existing text-based and TS-based methods.
@@
  • Comprehensive evaluations of VisionTS demonstrate its significant zero-shot performance, which even surpasses few-shot text-based TSF foundation models and achieves comparable or superior results to zero-shot TS-based models.
@@ 3 Methodology
As noted in the Introduction, TS and images share intrinsic similarities, suggesting the transfer potential of visual pre-trained models (particularly MAE in this paper) for TSF tasks. This section explains how to reformulate TSF tasks into MAE’s pre-training task, i.e., patch-level image reconstruction.
Our idea is straightforward: map the look-back/forecasting windows to visible/masked patches, respectively. This idea is supported by the recent success of prompt tuning (Schick & Schütze, 2021) in NLP, where the predictions for [mask] token in pre-trained language models (e.g., BERT (Devlin et al., 2019)) are directly used for downstream tasks. By unifying the forms of the two tasks, we bridge the gap between the two domains, enabling a MAE for zero-shot TSF directly without adapting the pre-trained parameters.
Notably, this idea is limited to univariate forecasting since multivariates are intractable to be encoded in a single image. Fortunately, recent work shows that channel independence — predicting each variable separately for multivariate forecasting — can still be effective (Nie et al., 2022; Han et al., 2024). Therefore, we leave the exploration of multivariate interactions for future work.
@@
[…]
@@
MAE standardizes each image based on the mean and standard deviation computed on ImageNet. Therefore, we apply instance normalization to ${\bm{I}}_{\text{raw}}$, which is also a standard practice in current TSF (Kim et al., 2022). Notably, we observed that normalizing ${\bm{I}}_{\text{raw}}$ to a standard deviation of $r$, where $r$ is a hyperparameter less than 1, yields superior performance. One explanation is that the magnitude of inputs/outputs during MAE pretraining is constrained by the limited range of color values. Therefore, reducing the magnitude of ${\bm{I}}_{\text{raw}}$ prevents exceeding these limits. However, an excessively low $r$ can result in values that are difficult to distinguish. We found that a moderate value (0.4) of $r$ performs well across most scenarios (Section B.9). Let ${\bm{I}}_{\text{norm}}$ denote the normalized matrix, which is computed as follows:
@@
It is well-known that each image has three channels. We simply render ${\bm{I}}_{\text{norm}}$ as a grayscale image ${\bm{I}}_{\text{grey}}\in\mathbb{R}^{P\times\lfloor\nicefrac{{L}}{{P}}\rfloor\times 3}$, where all three channels are identical to ${\bm{I}}_{\text{norm}}$. This choice is purely result-driven: In our early experiments, we added a convolutional layer with three output channels to convert the grayscale image into a color image and then fine-tuned it to find the optimal color transformation, which, however, did not significantly influence the performance.
@@
Moreover, we found that reducing the width of the visible portion can further improve performance, likely because the number of input periods ($\lfloor\nicefrac{{L}}{{P}}\rfloor$) is generally smaller than the period length $P$. Therefore, we propose multiplying $n$ by a hyperparameter $c\in[0,1]$ to reduce the width, i.e.,
$n=\left\lfloor c\cdot N\cdot\frac{L}{L+H}\right\rfloor.$
@@
4 Experiments
We use MAE (Base) with 112M parameters as our backbone. We conduct experiments on popular TSF benchmarks, including long-term TSF datasets (Zhou et al., 2021; Wu et al., 2021) and Monash benchmark (Godahewa et al., 2021). We select representative baselines for comparison, including TS-based foundation models: Moirai (Woo et al., 2024) and TimesFM (Das et al., 2024); Text-based foundation models: TimeLLM (Jin et al., 2024), GPT4TS (Zhou et al., 2023), and LLMTime (Gruver et al., 2023); and other popular TSF baselines covering both Transformer-based, MLP-based and CNN-based architectures. Details are elaborated in Appendix A.
4.1 Zero-Shot Time Series Forecasting
Setups
We first evaluate VisionTS’s zero-shot TSF performance without any fine-tuning on time-series modalities. To prevent data leakage and assess the out-of-distribution capabilities, we selected six widely-used datasets from the long-term TSF benchmark that are not included in Moirai’s pre-training set for evaluation. Since most baselines cannot perform zero-shot forecasting, we report their few-shot results by fine-tuning on the 10% of the individual target datasets. We also evaluate the Monash benchmark including 29 test datasets, which is more challenging for VisionTS since they were used in Moirai’s pre-training but not for VisionTS. We set the hyperparameters to $r=c=0.4$. Following common practice (Nie et al., 2022; Zhou et al., 2023; Woo et al., 2024), we conduct hyperparameter tuning on validation sets to determine the optimal context length $L$, detailed in Section B.1.
Table 1: Zero-shot or few-shot results on the long-term TSF benchmark. Results are averaged across prediction lengths {96, 192, 336, 720}, with full results in Section B.2. Bold: the best result. Comparisons with TimesFM and LLMTime are detailed in Section B.3.
@@ 4.1 Zero-Shot Time Series Forecasting
Results on Long-Term TSF Benchmark
Table 1 shows that VisionTS surprisingly achieves the best forecasting performance in most cases (7 out of 14). Specifically, VisionTS demonstrates a relative average MSE reduction of approximately 6% compared to MoiraiSmall and MoiraiLarge, and performs comparably to MoiraiBase. When compared to the various few-shot baselines, VisionTS shows a relative average MSE reduction ranging from 8% to 84%. Given that all baselines except for VisionTS are trained on the time-series domain, this result is particularly encouraging. It suggests that the transferability from images to time-series is stronger than from text to time-series, and even comparable to the in-domain transferability between time-series. We also include a comparison with traditional algorithms (ETS, ARIMA, and Seasonal Naïve) in Section B.4, where VisionTS still outperforms all of these traditional methods.
Results on Monash Benchmark
Fig. 4 shows the results aggregated from 29 Monash datasets, showing that VisionTS in the zero-shot setting surpasses all models individually trained on each dataset (e.g., FFNN, WaveNet, and TBATS) and significantly outperforms the other cross-domain baseline (i.e., LLMTime). It achieves second place among all baselines, just behind Moirai that pre-trained on all the training datasets. This highlights VisionTS’s strong zero-shot forecasting ability and the effective cross-modality transferability between image and time series.
In Appendix D, we visualize the zero-shot forecasting of VisionTS alongside the input and reconstruction images, highlighting both successful cases (where VisionTS outperforms Moirai) and failures (where Moirai prevails). When the input exhibits strong regularity (Fig. 8), VisionTS effectively forecasts both the periodicity (via segmentation) and trends (via MAE’s capabilities). In contrast, Moirai, akin to seasonal naïve methods, struggles to capture inter-period trends. For less-structured input (Figs. 9, 10 and 11), Moirai adopts a conservative approach with lower volatility to minimize errors, while VisionTS takes a more aggressive stance. This strategy occasionally yields more accurate trend predictions (Figs. 9 and 10) but may also result in greater MAE (Fig. 11).
Figure 4: Aggregated results on the Monash TSF Benchmark, with full results in Section B.5.
@@
Avg. 0.345 0.346 0.349
Table 2: MSE of different MAE variants, averaged on four prediction lengths.
Analysis: Impact of MAE’s parameter size

### v3

@@ Abstract
Foundation models have emerged as a promising approach in time series forecasting (TSF). Existing approaches either repurpose large language models (LLMs) or build large-scale time series datasets to develop TSF foundation models for universal forecasting. However, these methods face challenges due to the severe cross-domain gap or in-domain heterogeneity. This paper explores a new road to building a TSF foundation model from rich, high-quality natural images. Our key insight is that a visual masked autoencoder, pre-trained on the ImageNet dataset, can naturally be a numeric series forecaster. By reformulating TSF as an image reconstruction task, we bridge the gap between image pre-training and TSF downstream tasks. Surprisingly, without further adaptation in the time series domain, the proposed VisionTS could achieve better zero-shot forecast performance than existing TSF foundation models. With fine-tuning for one epoch, VisionTS could further improve the forecasting and achieve state-of-the-art performance in most cases. Extensive experiments reveal intrinsic similarities between images and real-world time series, suggesting that visual models may offer a “free lunch” for TSF and highlight the potential for future cross-modality research. Our code is available in the Supplementary Material.
Time Series Forecasting, Foundation Models, Transfer Learning
Figure 1: Long-term forecasting performance comparison. Our VisionTS, without any training on time series data, outperforms the pure time series foundation model MoiraiLarge in the zero-shot setting.
@@ 1 Introduction
We have recently witnessed two roads to building a TSF foundation model for universal forecasting. The first tries to repurpose large language models (LLMs) that have been pre-trained on text data for TSF tasks (i.e., text-based) (Zhou et al., 2023; Jin et al., 2024), based on the observation that LLMs and TSF models share a similar left-to-right forecasting paradigm. However, due to the significant gap between these two modalities, the effectiveness of such transferability between language and time series has recently been questioned by Tan et al. (2024).
@@
The second road focuses on constructing large-scale time-series datasets collected from diverse domains to train a TSF foundation model from scratch (i.e., time series-based or TS-based) (Woo et al., 2024; Das et al., 2024). Nevertheless, unlike images or language with unified formats, time series data is highly heterogeneous in length, frequency, number of variates, domains, and semantics, limiting the transferability between pre-training and downstream domains. Until recently, constructing a high-quality dataset remains challenging and is still in the early exploration stage.
@@
In this paper, we investigate a third road that is less explored yet promising: building TSF foundation models with pre-trained visual models. Our key idea is that pixel variations in a natural image can be interpreted as temporal sequences, which share many intrinsic similarities with time series: 1 Similar modalities: Unlike discrete texts, both images and time series are continuous; 2 Similar origin: Both time series and images are observations of real-world physical systems, whereas languages are products of human cognitive processes; 3 Similar information density: Languages are human-generated signals with high semantic density, while images and time series are natural signals with heavy redundancy (He et al., 2022); and 4 Similar features: As shown in Fig. 2, images often display many features of real-world time series, which are rarely found in language data. Based on these findings, images could be a promising modality for transferring to TSF. We are motivated to answer the question: Can a visual model pre-trained on images be a free-lunch foundation model for time series forecasting?
@@
Figure 2: An image of the ImageNet dataset (Deng et al., 2009), in which the pixel arrays can display many well-known features of real-world time series, such as trend, seasonality, and stationarity (Qiu et al., 2024). By self-supervised pre-training on ImageNet, it is reasonable that a visual model could understand these features and exhibit a level of time series forecasting ability.
@@
We focus on visual masked autoencoder (MAE) [^1] We use fonts to distinguish MAE (Masked Autoencoder) and MAE (Mean Absolute Error) in this paper., a popular CV foundation model (He et al., 2022) by self-supervised pre-training on ImageNet (Deng et al., 2009). As an image reconstruction and completion model, MAE can naturally be a numeric series forecaster. Inspired by the well-known prompt technique in NLP (Schick & Schütze, 2021), we propose a simple method to reformulate TSF as a patch-level image reconstruction task to bridge the gap between pre-training and downstream tasks. Specifically, we transform 1D time-series data into 2D matrices via segmentation. Then, we render the matrices into images and align the forecasting window with masked image patches. This allows us to make a zero-shot forecast without further adaptation.
@@
We evaluate our proposed VisionTS on large-scale benchmarks, including 8 long-term TSF (Zhou et al., 2021), 29 Monash (Godahewa et al., 2021), and 23 GIFT-Eval (Aksu et al., 2024) datasets, spanning diverse domains, frequencies, and multivariates. To the best of our knowledge, the scale of our evaluation benchmark is the largest among existing TSF foundation models. As demonstrated in Fig. 1, without further adaptation on time series, a vanilla MAE can surprisingly achieve a comparable performance or even outperform the strong zero-shot TSF foundation models. By fine-tuning MAE in each downstream dataset for a single epoch, VisionTS can lead to SOTA performance in most long-term TSF benchmarks.
To further understand and explain the transferability, we use an MAE encoder to visualize both modalities, showing a level of similarity between time series and natural image representations. Additionally, we observe considerable heterogeneity within time-series data across domains, and images can serve as a bridge to connect these isolated time-series representations. This could further explain why VisionTS performs better than some cross-domain TSF models. Our findings suggest that time series and natural images may be two sides of a coin, and visual models can be a free lunch for time series forecasting. We hope our findings inspire future cross-modal research on CV and TSF.
@@
  • We explore a road to building a TSF foundation model from natural images, conceptually different from the existing text-based and TS-based pre-training methods.
@@
  • Comprehensive evaluations of VisionTS on large-scale benchmarks across multiple domains demonstrate its significant forecasting performance, surpassing few-shot text-based TSF foundation models and achieving comparable or superior results to zero-shot TS-based models.
@@ 3 Methodology
As noted in the Introduction, time series and images share intrinsic similarities, suggesting the transfer potential of pre-trained visual models (particularly MAE in this paper) for TSF. To reformulate TSF tasks into MAE’s pre-training task, our high-level idea is straightforward: map the look-back/forecasting windows to visible/masked patches, respectively. This idea is supported by the prompt tuning (Schick & Schütze, 2021) in NLP, where the predictions for [mask] token in pre-trained language models, e.g., BERT (Devlin et al., 2019), are directly used for downstream tasks. By unifying the forms of the two tasks, we bridge the gap between the two modalities without further training.
@@
Given a univariate input $X\in\mathbb{R}^{L}$, the first goal is to transform it into a 2D matrix. We propose to segment it into $\lfloor\nicefrac{{L}}{{P}}\rfloor$ subsequences of length $P$, where $P$ is the periodicity. Notably, when the time series lacks clear periodicity, we can set $P=1$ directly, which is also effective in our experiments (Section C.5). In practice, $P$ can be determined using statistical methods like Fast Fourier Transform (Wu et al., 2023; Chen et al., 2024) or domain knowledge like sampling frequency (Godahewa et al., 2021; Alexandrov et al., 2020). In this paper, we select $P$ based on the sampling frequency, elaborated in Section B.2.
After that, these subsequences are then stacked into a 2D matrix, denoted by ${\bm{I}}_{\text{raw}}\in\mathbb{R}^{P\times\lfloor\nicefrac{{L}}{{P}}\rfloor}$. This encoding strategy is proven to be efficient by recent work like TimesNet (Wu et al., 2023) and SparseTSF (Lin et al., 2024), as it allows for the simultaneous capture of both variations within the same period (i.e., intra-period) and across periods with the same phase (i.e., inter-period). Moreover, it ensures that each element in ${\bm{I}}_{\text{raw}}$ and its neighbors align with the spatial locality property of images (Krizhevsky et al., 2012), where nearby pixels tend to be similar due to the inherent cohesiveness of objects in the real world. Therefore, this further narrows the gap between time series and images.
@@
MAE standardizes each image based on the mean and standard deviation computed on ImageNet. Therefore, we apply instance normalization to ${\bm{I}}_{\text{raw}}$, which is also a standard practice in current TSF (Kim et al., 2022). Notably, we observed that normalizing ${\bm{I}}_{\text{raw}}$ to a standard deviation of $r$, where $r$ is a hyperparameter less than 1, yields superior performance. One explanation is that the magnitude of inputs/outputs during MAE pretraining is constrained by the limited range of color values. Therefore, reducing the magnitude of ${\bm{I}}_{\text{raw}}$ prevents exceeding these limits. However, an excessively low $r$ can result in values that are difficult to distinguish. We found that a moderate value (0.4) of $r$ performs well across most scenarios (See Section C.8 for more details). Let ${\bm{I}}_{\text{norm}}$ denote the normalized matrix, which is computed as follows:
@@
Since each image has three channels, we simply render ${\bm{I}}_{\text{norm}}$ as a grayscale image ${\bm{I}}_{\text{grey}}\in\mathbb{R}^{P\times\lfloor\nicefrac{{L}}{{P}}\rfloor\times 3}$, where all three channels are identical to ${\bm{I}}_{\text{norm}}$. This choice is purely result-driven: In our early experiments, we added a convolutional layer with three output channels to convert the grayscale image into a color image and then fine-tuned it to find the optimal color transformation, which, however, did not significantly improve the performance.
@@
Moreover, we found that reducing the width of the visible portion can further improve performance. One possible explanation is that MAE uses a large masked ratio during pre-training, with only 25% of patches visible. Reducing the image width may align the masked ratio more closely with pre-training. Therefore, we propose multiplying $n$ by a hyperparameter $c\in[0,1]$. Similar to $r$, we found that setting $c=0.4$ performs well in our experiments (See Section C.8). This can be formulated as $n=\left\lfloor c\cdot N\cdot\nicefrac{{L}}{{(L+H)}}\right\rfloor$.
@@
Table 1: Zero-shot or few-shot results on the long-term TSF benchmark. Results are averaged across prediction lengths {96, 192, 336, 720}, with full results in Table 9 (Section C.2). Bold: the best result.
@@
Pretrain [Uncaptioned image] Images [Uncaptioned image] Time series [Uncaptioned image] Text [Uncaptioned image] No Pretrain
@@
Discussion on Multivariate Forecasting
In addition to the temporal interactions, multivariate time series data sometimes show interactions between variables. While pre-trained vision models effectively capture temporal interactions based on the intrinsic similarities between images and time series, they struggle to capture inter-variable interactions due to a limited number of image channels, especially without further training. Fortunately, recent work shows that channel independence — forecasting each variable separately — can be effective and is widely used in recent deep forecasting models (Nie et al., 2022; Han et al., 2024; Jin et al., 2024; Zhou et al., 2023; Lin et al., 2024). Following these works, we adopt channel independence in our paper while leaving the exploration of capturing inter-variable interactions to future work.
4 Experiments
Figure 4: Performance on the GIFT-Eval Leaderboard.
Figure 5: Aggregated results on the Monash TSF Benchmark, with full results in Table 13 (Section C.5).
We follow the standard evaluation protocol proposed by Woo et al. (2024) to test our VisionTS on 35 widely-used TSF benchmarks, and additionally evaluate it on the GIFT-Eval (Aksu et al., 2024) which is the largest TSF benchmark for zero-shot foundation models. We use MAE (Base) as our backbone by default. Baseline and benchmark details are elaborated in Section B.1.
4.1 Zero-Shot Time Series Forecasting
Setups
We first evaluate VisionTS’s zero-shot TSF performance without fine-tuning on time-series modalities. To prevent data leakage, we selected six widely-used datasets from the long-term TSF benchmark that are not included in Moirai’s pre-training set for evaluation. Since most baselines cannot perform zero-shot forecasting, we report their few-shot results by fine-tuning on the 10% of the individual target datasets. We also evaluate the Monash benchmark and GIFT-Eval benchmark. Notably, the Monash benchmark is more challenging for VisionTS since they were used in Moirai’s pre-training but not for VisionTS. We set the hyperparameters to $r=c=0.4$. Following common practice (Zhou et al., 2023; Woo et al., 2024), we conduct hyperparameter tuning on validation sets to determine the optimal context length $L$, detailed in Section C.1.
Table 2: Average MSE of different MAE variants, with full results in Table 15 (Section C.6).
@@ 4.1 Zero-Shot Time Series Forecasting
Electricity 0.207 0.209 0.202
Weather 0.269 0.272 0.292
Avg. 0.309 0.311 0.315
Context Length 1k 1k 2k 3k 4k
Prediction Length 1k 2k 3k 4k 1k
PatchTST 0.01 0.01 0.01 0.01 0.01 0.02 0.03 0.04
DeepAR 0.26 0.32 0.37 0.43 0.26 4.06 6.10 8.17
GPT4TS 0.01 0.01 0.01 0.02 0.01 0.03 0.04 0.06
MoiraiBase 0.03 0.04 0.04 0.05 0.03 0.04 0.05 0.06
TimesFM 0.08 0.14 0.20 0.27 0.07 0.13 0.20 0.25
LLMTime (8B) > 200 > 200
VisionTS ($c=0.4$) 0.04 0.03 0.03 0.03 0.04 0.04 0.05 0.05
Results on Long-Term TSF Benchmark
Table 1 shows that VisionTS surprisingly achieves the best forecasting performance in most cases (7 out of 14). Specifically, VisionTS demonstrates a relative average MSE reduction of approximately 6% compared to MoiraiSmall and MoiraiLarge, and performs comparably to MoiraiBase. When compared to the various few-shot baselines, VisionTS shows a relative average MSE reduction ranging from 8% to 84%. Given that all baselines except for VisionTS are trained on the time-series domain, this result is particularly encouraging. It suggests that the transferability from images to time-series is stronger than from text to time-series, and even comparable to the in-domain transferability between time-series. We also include a comparison with two TSF foundation models, TimesFM (Das et al., 2024) and LLMTime (Gruver et al., 2023), in Section C.3, as well as traditional algorithms (ETS, ARIMA, and Seasonal Naïve) in Section C.4. Results show that VisionTS still outperforms all of these baselines.
Results on GIFT-Eval and Monash Benchmarks
Fig. 4 shows the comparison of VisionTS with six previously published TSF foundation models on the GIFT-Eval TSF Leaderboard [^2] https://huggingface.co/spaces/Salesforce/GIFT-Eval, where VisionTS surprisingly ranked first in terms of normalized MASE. Besides, Fig. 5 shows the results aggregated from 29 Monash datasets, showing that VisionTS in the zero-shot setting surpasses all models individually trained on each dataset and significantly outperforms the other cross-domain baseline (i.e., LLMTime). It achieves second place among all baselines, just behind Moirai that pre-trained on all the training datasets. This promising result highlights VisionTS’s strong zero-shot forecasting ability and effective cross-modality transferability.
4.2 Further Analysis of VisionTS
Backbone Analysis